### v1

# Solving Truly Massive Budgeted Monotonic POMDPs with Oracle-Guided Meta-Reinforcement Learning

Abstract
Monotonic Partially Observable Markov Decision Processes (POMDPs), where the system state progressively decreases until a restorative action is performed, can be used to model sequential repair problems effectively. This paper considers the problem of solving budget-constrained multi-component monotonic POMDPs, where a finite budget limits the maximal number of restorative actions. For a large number of components, solving such a POMDP using current methods is computationally intractable due to the exponential growth in the state space with an increasing number of components. To address this challenge, we propose a two-step approach. Since the individual components of a budget-constrained multi-component monotonic POMDP are only connected via the shared budget, we first approximate the optimal budget allocation among these components using an approximation of each component POMDP’s optimal value function which is obtained through a random forest model. Subsequently, we introduce an oracle-guided meta-trained Proximal Policy Optimization (PPO) algorithm to solve each of the independent budget-constrained single-component monotonic POMDPs. The oracle policy is obtained by performing value iteration on the corresponding monotonic Markov Decision Process (MDP). This two-step method provides scalability in solving truly massive multi-component monotonic POMDPs. To demonstrate the efficacy of our approach, we consider a real-world maintenance scenario that involves inspection and repair of an administrative building by a team of agents within a maintenance budget. Our results show that the proposed method significantly improves average component survival times compared to baseline policies, thereby highlighting its potential for practical applications in large-scale maintenance problems. Finally, we perform a computational complexity analysis for a varying number of components to show the scalability of the proposed approach.
1 Introduction
Partially Observable Markov Decision Processes (POMDPs) provide an efficient framework for modeling a wide range of real-world sequential decision-making problems (Cassandra 1998; Bravo, Leiras, and Cyrino Oliveira 2019). Numerous methods have been developed to compute optimal policies for POMDPs, including Monte-Carlo
Figure 1: Condition of infrastructure component over time. (a) Line plot showing component condition over time for 100 runs. The red x marks denote the time step when condition reaches 0. (b) Violin-plot showing distribution of component condition for different time steps.
methods (Katt, Oliehoek, and Amato 2017), reinforcement learning approaches (Singh et al. 2021), and various approximation techniques (Kearns, Mansour, and Ng 1999). One notable application of POMDPs is in infrastructure maintenance or sequential repair problems, where the system state represents the condition of infrastructure components, which is only partially observable (Bhattacharya et al. 2021). Over time, the state of these components stochastically deteriorates unless a restorative action, such as a repair, is performed. Figure 1 shows this stochastic decrease and also shows the probability distribution of the condition of a sample infrastructure component for different time steps. The state evolution of the POMDP, thus, exhibits monotonic properties (Miehling and Teneketzis 2020) and we define such POMDPs as monotonic POMDPs. Previously, Bhattacharya et al. (2020) proposed an optimal inspection and repair scheduling policy for a pipeline, which is a single-component POMDP. However, infrastructure systems naturally have multiple components (Daulat et al. 2024).
In this paper, we address the challenge of computing approximately optimal policies for budget-constrained multi-component monotonic POMDPs. Each component POMDP operates independently in terms of transition probabilities, but they are collectively constrained by the shared budget. There has been substantial work on solving budget constrained POMDPs (Lee et al. 2018; Undurti and How 2010; Khonji, Jasour, and Williams 2019). However, the complexity of these algorithms is often exponential in the number of states of the POMDP, and consequently would be exponential in the number of components for a multi-component POMDP. Thus, they become computationally intractable for multi-component POMDPs with a large number of components. In Vora et al. (2023), the authors propose a welfare-maximization method for solving budget-constrained multi-component POMDPs. However, the method in that paper requires generating optimal policies for multiple budget values for every component POMDP to get the optimal budget allocation, i.e., optimal distribution of the shared budget among the components POMDPs. Hence, it cannot be scaled to a large number of components.
To address this scalability issue, we propose a computationally efficient meta-reinforcement learning approach for solving budget-constrained multi-component monotonic POMDPs. First, we approximate the optimal distribution of budget among the individual component POMDPs using a random forest budget allocation strategy, thereby decoupling the problem into independent components. The random forest model is trained to approximate the value function of a component POMDP as a function of the budget, leveraging the concavity of the value function to frame the budget allocation as a concave maximization problem, which guarantees a computationally feasible solution. Subsequently, we apply the oracle policy-guided meta-trained Proximal Policy Optimization (PPO) algorithm, trained across various components and budget values, to compute the approximately optimal policy for each component-budget pair. The oracle policy for the meta-PPO is derived from value iteration applied to the corresponding component MDP. To validate the efficacy of our proposed approach, we examine a maintenance scenario involving an administrative building with 1000 infrastructure components and a fixed maintenance budget. We compare the performance of our approach, in terms of the survival time of the components, against the oracle policy, a currently-used baseline policy, and a vanilla meta-PPO algorithm. We also perform a computational complexity analysis of our algorithm for varying number of components and show that the proposed approach is linear in the number of components.
In summary, the key contributions of this paper are:
  1. 1.
    We develop an oracle policy-guided meta-PPO algorithm for solving budget-constrained single-component monotonic POMDPs.
  2. 2.
    We introduce a random forest budget allocation strategy for optimally distributing the budget among the component POMDPs of a budget-constrained multi-component POMDP.
  3. 3.
    We demonstrate the efficacy of our approach through a large-scale real-world infrastructure maintenance scenario with 1000 infrastructure components.
  4. 4.
    We show the scalability of our algorithm through a computational complexity analysis for varying number of components.
2 Preliminaries and Related Work
2.1 Partially Observable Markov Decision Processes
A discrete-time finite-horizon Partially Observable Markov Decision Process (POMDP) (Cassandra, Kaelbling, and Littman 1994) $M$ is defined by the 7-tuple $(\mathcal{S},A,T,\Omega,O,R,H)$, which denotes the state space, action space, state transition function, observation space, observation function, reward function and planning horizon, respectively. In a POMDP, the agent does not have direct access to the true state of the environment. Instead, the agent may maintain a belief state, representing a probability distribution over $\mathcal{S}$. This belief is updated based on the received observation using Bayes’ rule (Araya et al. 2010).
2.2 POMDP Solution Methods
Computing optimal policies for a POMDP is generally PSPACE-complete (Mundhenk et al. 2000; Vlassis, Littman, and Barber 2012). Thus, to address the computational intractability of solving POMDPs, various approximation methods have been widely used (Poupart and Boutilier 2002; Pineau et al. 2003; Roy, Gordon, and Thrun 2005). Several reinforcement learning approaches have also been developed for computing approximate POMDP solutions (Azizzadenesheli, Lazaric, and Anandkumar 2016; Igl et al. 2018). However these methods become computationally intractable when faced with the high dimensionality and shared resource constraints of budget-constrained multi-component monotonic POMDPs such as those considered in this paper.
2.3 Consumption MDPs and Budgeted POMDPs
The integration of budget or resource constraints into Markov Decision Processes (MDPs) has been previously studied under the frameworks of Consumption MDPs (Blahoudek et al. 2020) and Budgeted POMDPs (Vora et al. 2023). However, the algorithm proposed in Blahoudek et al. (2020) assumes full observability of the state and hence cannot be applied to budget-constrained POMDPs. A solution for budget-constrained multi-component POMDPs is presented in Vora et al. (2023). However, the method in this paper requires repeated computations of optimal policies for different budget values for all component POMDPs and hence is not scalable to a budget-constrained multi-component POMDP with a large number of components.
Figure 2: Architectural overview of the proposed approach.
3 Problem Formulation
In this paper, we consider a weakly-coupled multi-component monotonic POMDP with a total budget. A weakly-coupled multi-component POMDP refers to a system where the individual component POMDPs have independent transition probabilities but are interconnected through a shared budget $B$. This shared budget introduces a weak coupling between the components, as the allocation of budget to one component affects the available budget for the others. The state space for an $n$-component monotonic budget-constrained POMDP is given by $\mathcal{S}=\prod_{i=1}^{n}\mathcal{S}_{i},$ where $\mathcal{S}_{i}$ represents the state space for component $i$, and $i\in\{1,\ldots,n\}$. The action space is given by $\mathcal{A}=\prod_{i=1}^{n}\mathcal{A}_{i},$ where the action space for component $i$ is $\mathcal{A}_{i}=\{d^{i},q^{i},m^{i}\}$. Each action incurs a fixed cost. The state at time instant $k$ is an $n$-tuple, $s_{k}=(s_{k}^{1},s_{k}^{2},\cdots,s_{k}^{n}),$ where $s_{k}^{i}\in\mathcal{S}_{i}=\{0,1,\ldots,\bar{s}\}$ denotes the state of component $i$, and $\bar{s}\in\mathbb{N}_{0}$ is the maximum possible value of $s_{k}^{i}$. Here, $\mathbb{N}_{0}$ denotes the set of non-negative integers. Similarly, the action at time $k$ is given by $a_{k}=(a_{k}^{1},a_{k}^{2},\cdots,a_{k}^{n})$ and the cost associated with this action is given by $c_{a_{k}}=\sum_{i=1}^{n}c_{a^{i}_{k}},$ where $c_{a^{i}_{k}}$ represents the cost associated with each action $a_{k}^{i}$. The transition function for the multi-component POMDP is:
$T(s_{k},a_{k},s_{k+1})=\prod_{i=1}^{n}T_{i}(s^{i}_{k},a^{i}_{k},s^{i}_{k+1}).$
The transition probability function for each component $i$ is:
$T^{i}(s^{i}_{k},a^{i}_{k},s^{i}_{k+1})=\begin{cases}p_{1}^{i}(s^{i}_{k},a^{i}_{k},s^{i}_{k+1}),&\text{if }a^{i}_{k}=m^{i}\text{ and }\\ &s^{i}_{k}\leq s^{i}_{k+1}\leq\bar{s},\\ p_{2}^{i}(s^{i}_{k},a^{i}_{k},s^{i}_{k+1}),&\text{if }a^{i}_{k}\in\{d^{i},q^{i}\}\\ &\text{and }s^{i}_{k+1}\leq s^{i}_{k},\\ 1,&\text{if }a^{i}_{k}\in A^{i}\text{ and }\\ &s^{i}_{k+1}=0=s^{i}_{k},\\ 0,&\text{otherwise}.\end{cases}$
Here, action $m^{i}$ is a restorative action that increases the state value, with the increase being upper bounded by $\bar{s}$. In contrast, actions $d^{i}$ and $q^{i}$ decrease the state value. Moreover, $s_{k}^{i}=0$ is an absorbing state for all $k,i$. Finally, the observation probability function for each component follows the model from Vora et al. (2023), where action $q_{i}$ gives true state information and the other two actions provide no information about the true state.
3.1 Problem Statement
The primary objective of this paper is to determine a policy $\pi^{*}$ for this multi-component monotonic POMDP over a horizon $H$, that maximizes the sum of expectations of the individual times before reaching the absorbing state for each component, while adhering to the total budget $B$. We denote this maximal time $k$ by $T_{max}=\sum_{i=1}^{n}T^{i}_{max}$, where $T^{i}_{max}$ denotes the corresponding maximal time for component $i$. Mathematically, the problem can be formulated as:
$\max_{\pi}\left(\sum_{i=1}^{n}\mathbb{E}[T^{i}_{max}(\pi)]\right)$ (1)
[…]
In this formulation, $\pi$ represents the policy, and both $T^{i}_{max}$ and $c_{a_{k}}$ depend on $\pi$. For simplicity, we will not explicitly denote this dependence in the remainder of this paper. There are many other possible formulations of the objective of the problem statement like a maxmin formulation:
$\max_{\pi}\min_{i}\mathbb{E}[T^{i}_{max}(\pi)].$ (2)
In this paper we consider the formulation given by (1).
4 Solution Approach
In this section, we present our methodology for solving a budget-constrained multi-component monotonic POMDP. Figure 2 presents an architectural overview of our proposed approach. We first input information about all $n$ component POMDPs into a pre-trained random forest regressor to get the $T^{i}_{max}$ for each component $i$ as a function of the budget allocated to the component. Next, we propose an appropriate allocation of the shared budget $B$ among the component POMDPs by solving a constrained maximization problem. We then compute the oracle policy for each component POMDP and budget pair through value iteration applied to the corresponding component MDP. Finally, using these oracle policies and a meta-trained reinforcement learning (RL) agent, we solve each component POMDP with respect to the allocated budget and consequently propose a policy for the multi-component POMDP. Note that an alternate allocation strategy could involve redistributing the budget at every time step during planning. However, such a method would be computationally more expensive, due to the repeated computation of the allocation, as compared to our proposed a priori budget distribution approach.
4.1 Oracle-Guided RL for Budget-Constrained Monotonic POMDPs
To derive the optimal policy for a budget-constrained single-component monotonic POMDP, we propose an oracle policy-guided reinforcement learning algorithm. Adherence to the budget constraints is achieved using the budget-constrained POMDP (bPOMDP) framework proposed in Vora et al. (2023). In a bPOMDP, the state at each time step includes an additional fully observable component representing the cost incurred up to that point.
The oracle policy is denoted as $\pi_{\text{oracle}}$ and is obtained by solving the corresponding MDP using value iteration. For a single-component monotonic POMDP with budget $B$, the corresponding MDP has an action space $\mathcal{A}_{\text{MDP}}=\{d,m\}$, identical transition probabilities as the POMDP, and full observability of the state. We use the Proximal Policy Optimization (PPO) algorithm (Schulman et al. 2017) in conjunction with the oracle policy to solve the single-component budget-constrained POMDP. Doing so allows us to treat the problem as that of determining the optimal inspection policy. At each time step, the PPO agent decides whether to perform inspection or not. If the agent decides to take the inspection action, it receives accurate information about the true state of the POMDP. On the other hand, if the agent decides not to inspect, then the action is chosen according to the oracle policy. Hence, the action space for the PPO agent is reduced to $\{q,\neg q\}$. Since the full state is not observable in a POMDP, we utilize the belief $b_{s}$ for planning. The agent’s belief of the true state is updated at each time step using a particle filter approach. For our work, we empirically observe that using the expected belief $\bar{b_{s}}$ and the variance of the belief $\sigma^{2}_{b_{s}}$ suffices for planning.
Hence, for the proposed oracle policy-guided PPO agent, the state at time instant $k$ is given by the vector $[\bar{b}_{s_{k}},c_{k},\sigma^{2}_{b_{s_{k}}}]$. Furthermore, the reward function is defined as follows:
$R(s_{k},c_{k},a_{k})=\begin{cases}r_{1}<0,&\text{if }c_{k}>B,\\ r_{2}<0,&\text{if }\left\lfloor{\bar{b}_{s_{k}}}\right\rfloor=0,\\ r_{3}=\frac{k}{H}-\alpha|\bar{b}_{s_{k}}-s_{k}|,&\text{if }\bar{b}_{s_{k}},c_{k}>0,\end{cases}$
[…]
4.2 Random Forest Approach for Optimal Budget Allocation
In the previous subsection, we detailed the methodology for deriving the optimal policy for a single-component budget-constrained monotonic POMDP. We now extend this to an $n$-component POMDP described in Section 3. Each component in this multi-component POMDP operates independently in terms of transition probabilities, but they share a common total budget, rendering them weakly coupled. While reinforcement learning algorithms have made significant advances, they often face challenges when scaling to the extremely large state and action spaces characteristic of multi-component systems (Sutton and Barto 2018). In the case of a budget-constrained $n$-component POMDP, directly applying a single-component solution approach would result in an exponentially large state space and action space, making such an approach impractical for large-scale problems. To address this scalability issue, we propose an a priori budget distribution strategy to decouple the components, thereby reducing the size of the state and action spaces. Under certain assumptions, the expected maximal time to reach the absorbing state, $\mathbb{E}[T_{\max}]$, for a budget-constrained single-component monotonic POMDP, is concave with respect to the allocated budget $b$ (Vora et al. 2023). For a given component $i$ and allocated budget $b^{i}$, let $\tilde{T}^{i}_{\max}$ be the approximate expected maximal time as a function of the budget, modeled using an exponential function:
$\tilde{T}^{i}_{\max}=\alpha^{i}e^{\beta^{i}b^{i}}+\gamma^{i},$ (3)
where $\alpha^{i},\beta^{i},\gamma^{i}\in\mathbb{R}$ are constants. While many other concave functions could be used to model $\tilde{T}^{i}_{\max}$, we empirically observe that the exponential function provides a good fit for the data. We use a random forest regressor (Breiman 2001) to estimate the parameters of this exponential function. The training dataset for this model is obtained via non-linear least squares regression on multiple $(\mathbb{E}[T_{\max}],b)$ pairs for various budget-constrained single-component monotonic POMDPs. The input to this model includes specific statistics related to the POMDP’s transition function, which are the expected time to reach state 0 without repairs, $\mathbb{E}[T]$, and the variance of this expected time, $\sigma^{2}_{\mathbb{E}[T]}$, as well as the cost of the various actions.
For an $n$-component POMDP with total budget $B$, let the budget allocated to the $i$-th component be $b^{i}$, and the approximate expected maximal time to reach the absorbing state for this budget be $\tilde{T}^{i}_{\max}$.
Figure 3: Performance comparison of oracle policy, oracle-guided meta-PPO, realistic baseline and vanilla meta-PPO. (a) Comparison of $\hat{T}_{max}$ values for all 1000 components across different budget values allocated to each component. (b) Comparison of average number of repairs performed by the agent under each of the four policies. (c) Comparison of average total cost incurred by the agent over the planning horizon for each of the four policies.
We then formulate the following constrained maximization problem:
$\max_{b^{i}}\sum_{i=1}^{n}\tilde{T}^{i}_{\max}$ (4)
[…]
Solving (4) yields the approximately optimal budget allocation among the individual components. Since we assume $\tilde{T}^{i}_{\max}$ to be concave in $b^{i}$, the optimization problem becomes a concave maximization problem and is thus guaranteed to have a computationally tractable solution.
4.3 Optimal Policy for Multi-Component Monotonic POMDPs
We now integrate the approaches described in the previous subsections to compute the optimal policy for an $n$-component POMDP, where $n$ is substantially large. Utilizing the random forest regressor from Section 4.2, we efficiently approximate $\mathbb{E}[T_{\max}]$ for each component $i$. Additionally, we meta-train the oracle-guided PPO agent by constantly updating the policy network’s parameters over a randomly selectied a subset of components and budget values. This approach allows the agent to generalize across components. This meta-trained agent is then utilized to derive the optimal policy $\pi^{i^{*}}$ for each component $i$, following the optimal budget allocation obtained from (4). Consequently, the overall policy for the multi-component POMDP is:
$\pi^{*}(s_{k},a_{k})=(\pi^{1^{*}}(s_{k}^{1},a_{k}^{1}),\pi^{2^{*}}(s_{k}^{2},a_{k}^{2}),\cdots,\pi^{n^{*}}(s_{k}^{n},a_{k}^{n})).$
While this policy is not guaranteed to be globally optimal for the entire multi-component POMDP, we empirically observe that it performs well in practice while respecting the budget constraints. We validate this approach by evaluating its performance on real-world data in the subsequent section.
5 Implementation and Evaluation
In this section, we evaluate the efficacy of the proposed methodology for determining the optimal policy for a very large multi-component budget-constrained POMDP. Specifically, we compare our approach against existing methods in the context of a multi-component building maintenance scenario managed by a team of agents. We also perform a computational complexity analysis of the proposed approach, for varying number of components.
We consider an administrative building comprising 1000 infrastructure components, including roofing elements, water fountains, lighting systems, and boilers. Each component’s health is quantified by the Condition Index (CI) (Grussing, Uzarski, and Marrano 2006), which ranges from 0 to 100. For each infrastructure component, we utilize historical CI data to generate the transition probabilities for the corresponding POMDP, modeled using the Weibull distribution (Grussing, Uzarski, and Marrano 2006). We use the weibull_min class from the scipy.stats module in Python to simulate the CI transitions over time. While a seed can be set using the random_state parameter in weibull_min for reproducibility, we did not set one to preserve the stochastic nature of the CI transitions. The condition index deteriorates stochastically over time, influenced by various factors, and can only be accurately assessed through explicit inspections, which incur a cost. A component is considered to have failed when its CI falls below a failure threshold, which is assumed to be 0. Components can be repaired to increase their CI. The building is allocated a maintenance budget of $B=$ 500,000 units for a given horizon of 100 decision steps. At the beginning of the horizon, the CI of all components is 100. The objective of the agents is to maximize the time until failure of the components by efficiently allocating the budget among the components and performing repairs and inspections as needed. The replacement costs (ranging from 50 to 500 units) and inspection costs (ranging from 1 to 5 units) of these components are derived from industry averages. Consistent with the approach described in Section 4.1, we model this objective as a POMDP (with $\alpha=10^{-3}$ in the reward function). This POMDP has roughly $10^{2000}$ states and $3^{1000}$ actions.
5.1 Analysis of Maintenance Policy
We begin by analyzing the performance of the maintenance policy derived using the proposed oracle-guided meta-PPO strategy for a single-component POMDP representing a component $i$ of the 1000 components. This policy is compared with the performance of the oracle policy on the corresponding component MDP. Since the oracle policy has full observability of the state, it is expected to always perform better than the proposed approach. Additionally, we evaluate two baseline policies:
  1. 1.
    A heuristic policy often used in practice (Lam and Yeh 1994; Straub 2004) where the agent performs inspections at regular intervals and repairs the component when its expected belief about the Condition Index (CI) falls below a predefined threshold. We chose an inspection interval of 5 steps and a repair threshold of 15 after extensive experiments with intervals ranging from 1 to 10 steps and repair thresholds from 5 to 50.
  2. 2.
    A vanilla meta-PPO agent that is trained on the same subset of component-budget pairs as the oracle-guided agent, but without an oracle policy.
Both the oracle-guided meta-PPO and vanilla meta-PPO are trained for 2M time steps each, with an Adam stepsize of $10^{-4}$, a minibatch size 128, policy update horizon of $T=4096$ and discount factor $0.95$. All other hyperparameters follow those used in Schulman et al. (2017). We perform 100 simulations for this component to obtain the corresponding $T^{i}_{\max}$ values, which are then averaged over the runs for a given budget value allocated to the component. This process is repeated for all 1000 components and the run-averaged $T^{i}_{\max}$ values are then averaged across components. We compare this average denoted by $\hat{T}_{max}$ for 11 different budget values ranging from 0 to 5000 units, along with the average number of repairs performed by the agent and the average cost incurred over the planning horizon. Figure 3 illustrates a comparison of these metrics for all four policies. We observe that the proposed approach significantly outperforms the baselines. The oracle-guided meta-PPO agent nearly matches the performance of the oracle policy for all 3 metrics, presumably due to the low inspection costs of the components. If inspection costs were significantly higher, the agent’s performance would likely diverge from the oracle policy, which is an expected outcome given the budget constraints. We also infer that the vanilla meta-PPO agent has only learnt to not violate the budget constraint by not performing any repairs. These results demonstrate the value of incorporating an oracle policy into the training of a reinforcement learning agent.
5.2 Analysis of Budget Allocation
Next, we demonstrate the effectiveness of our random forest-based budget allocation strategy. We compare it with a baseline approach that allocates budgets proportional to the ratio of a component’s replacement cost to its $\mathbb{E}[T]$. For a component $i$, we model its $\mathbb{E}[T^{i}_{\max}]$ using $\tilde{T}^{i}_{\max}$ as given in (3). The parameters $\alpha^{i}$ and $\gamma^{i}$ can be estimated directly by considering the boundary conditions: $\gamma^{i}$ is estimated by substituting $b^{i}=0$, representing the scenario where no budget is available, and $\alpha^{i}$ is determined by substituting $b^{i}=\infty$, corresponding to the scenario of unlimited budget, where the supremum of $T^{i}_{\max}$ ($\sup_{b^{i}}T^{i}_{\max}=H=100$) is reached. We then train a random forest regressor to estimate parameter $\beta^{i}$. The training dataset is created by performing non-linear least squares regression on 11 distinct ($T^{i}_{max},b^{i}$) pairs each for 800 components. These pairs correspond to the run-averaged $T^{i}_{\max}$ values and the respective budget values $b^{i}$ from Section 5.1. The input to the random forest model is a vector consisting of the shape and scale factors of the Weibull distribution, which represent $\mathbb{E}[T]$ and $\sigma^{2}_{\mathbb{E}[T]}$, along with the replacement and inspection costs for a given component $i$. If a different distribution were used to model the transition probability, we would similarly extract the parameters, $\mathbb{E}[T]$ and $\sigma^{2}_{\mathbb{E}[T]}$, for inclusion in the input vector.
Figure 4: Performance of random forest model for predicting the value of parameter $\beta$ for a test dataset of 200 components. The horizontal axis represents parameter values obtained via non-linear least squares and vertical axis represents predicted values. The dotted line represents the $y=x$ line, i.e., perfect predictions.
Figure 4 shows the prediction performance of the random forest model for a test dataset of 200 components which were not encountered during training. We see that most points on the plot are very close to the perfect prediction line and bad predictions are few in number (29 out of 200 for error threshold of $10^{-4}$). The random forest model achieves a mean squared error (MSE) = $1.8\times 10^{-8}$ for this test dataset. Note that the non-linear least squares regressor constrains $\beta^{i}$ to be $\leq 0$ and hence for some components we observe that $\beta^{i}=0$. We use this trained random forest model to estimate $\tilde{T}^{i}_{max}$ for all 1000 components. Finally, using these approximated expressions, we solve the constrained maximization problem described in (4) to obtain the appropriate budget allocation for the components.
Approach $T_{max}$
Random Forest Budget Allocation 22,009.5
Baseline Budget Allocation 16,445.4
Table 1: $T_{max}$ (in steps), averaged over 100 runs, achieved under random forest and baseline budget allocations.
Figure 5: Performance comparison of random forest-based budget allocation and baseline budget allocation for all 1000 components for an overall budget of 500,000 units.
We quantify the performance of the random forest budget allocation and the baseline budget allocation algorithms by calculating the $T_{max}=\sum_{i}T^{i}_{max}$ and averaging it over 100 runs. For a fair comparison, these values are obtained using the oracle-guided meta-PPO approach for both allocation schemes.
Number of Components Random Forest Budget Split Value Iteration Meta-PPO
1 0.9724 0.9046 113.7227 2.8885
2 0.8870 0.8314 116.3281 3.0858
5 0.8719 0.8207 135.3953 4.7940
10 0.8762 0.8132 280.4909 9.5495
20 0.9534 0.8997 451.2948 16.2575
50 0.9449 0.8916 1208.1000 33.7387
100 0.9324 0.9171 2389.5641 64.6809
500 0.9575 1.2226 10269.1037 313.9742
1000 0.9599 1.6232 20612.1734 627.7477
Table 2: Time taken (in seconds) for running each process with varying numbers of components, averaged over 10 runs.
Table 1 shows the $T_{max}$ values achieved by both allocation approaches. The random forest budget allocation vastly outperforms the baseline approach. Furthermore, Figure 5 presents violin plots showing the distribution of the $T^{i}_{max}$ values achieved under the proposed and baseline budget allocations for all 1000 components. We observe that there are more components with higher $T^{i}_{max}$ values for the random forest budget allocation approach. Preliminary experiments on alternative objective formulations, such as the maxmin approach given by (2), also indicate that the proposed method consistently outperforms the baseline.
5.3 Analysis of Time Complexity
Finally, we analyze the time complexity of our proposed approach for varying number of components $N$. As mentioned earlier, our method comprises of four major steps:
  1. 1.
    Random Forest regression for estimating $\tilde{T}^{i}_{max}$ for each component $i$.
  2. 2.
    Budget Allocation among components via constrained optimization.
  3. 3.
    MDP Value Iteration for each component-budget pair to obtain the corresponding oracle policy.
  4. 4.
    Oracle-Guided Meta-PPO to approximately solve each component POMDP.
Table 2 presents the times taken for running each of the four processes, with different number of components. The time complexity experiments were performed in Python on a laptop running MacOS with an M2 chip @3.49GHz CPU and 8GB RAM. The times taken for random forest and budget allocation steps are negligible compared to those for performing value iteration and generating optimal policies through meta-PPO. The value iteration is applied to each component independently and hence scales linearly with the number of components. Similarly, Step 4 involves applying the pre-trained policy to each component separately and thus is also linear in the number of components. Consequently, we expect that the time complexity of our algorithm is linear in the number of components, i.e., $O(n)$. This expectation is confirmed by the log-log plot of computational complexity shown in Figure 6. Our algorithm’s performance is thus significantly faster as compared to existing POMDP solvers which would be exponential in the number of states and thus doubly exponential in the number of components (Silver and Veness 2010), (Pineau et al. 2003). If the problem is approached directly as a single POMDP, it will have a prohibitively vast state space of approximately $10^{2000}$ states. Previous work by Vora et al. (2023) demonstrated that standard methods indeed become computationally intractable after a few components due to this combinatorial explosion.
Figure 6: Log-log plot of computational complexity of the proposed approach for varying numbers of components.
6 Conclusions
In this paper we introduce a framework to provide computationally efficient solutions, scalable for large scale budget-constrained multi-component monotonic POMDPs. Noting that the coupling is expressed solely through the shared budget, our method functions by first decomposing the large $n$-component POMDP into $n$ independent single-component POMDPs using a random forest budget allocation of the shared budget. Next, we compute the oracle policy using value iteration for each component POMDP and budget pair. Finally, using this oracle policy together with a meta-trained PPO agent, we solve each component POMDP. Consequently, we obtain the approximately optimal policy for the multi-component POMDP. The experimental evaluations conducted for a real-world scenario of large scale infrastructure maintenance show that the proposed approach outperforms existing baselines and almost matches the performance of the oracle policy, which relies on information not available to the planner in a partially observable setting. Furthermore, computational complexity analysis demonstrates the scalability of our framework. Future work will focus on extending the framework’s capabilities to more dynamic budget allocation schemes and more complicated hierarchical budget constraints.

### v2

@@
 Abstract
-Monotonic Partially Observable Markov Decision Processes (POMDPs), where the system state progressively decreases until a restorative action is performed, can be used to model sequential repair problems effectively.
-This paper considers the problem of solving budget-constrained multi-component monotonic POMDPs, where a finite budget limits the maximal number of restorative actions.
-For a large number of components, solving such a POMDP using current methods is computationally intractable due to the exponential growth in the state space with an increasing number of components.
-To address this challenge, we propose a two-step approach.
-Since the individual components of a budget-constrained multi-component monotonic POMDP are only connected via the shared budget, we first approximate the optimal budget allocation among these components using an approximation of each component POMDP’s optimal value function which is obtained through a random forest model.
-Subsequently, we introduce an oracle-guided meta-trained Proximal Policy Optimization (PPO) algorithm to solve each of the independent budget-constrained single-component monotonic POMDPs.
-The oracle policy is obtained by performing value iteration on the corresponding monotonic Markov Decision Process (MDP).
-This two-step method provides scalability in solving truly massive multi-component monotonic POMDPs.
-To demonstrate the efficacy of our approach, we consider a real-world maintenance scenario that involves inspection and repair of an administrative building by a team of agents within a maintenance budget.
-Our results show that the proposed method significantly improves average component survival times compared to baseline policies, thereby highlighting its potential for practical applications in large-scale maintenance problems.
-Finally, we perform a computational complexity analysis for a varying number of components to show the scalability of the proposed approach.
+Many real-world decision problems, ranging from asset-maintenance scheduling to portfolio rebalancing, can be naturally modelled as budget-constrained multi-component monotonic Partially Observable Markov Decision Processes (POMDPs): each component’s latent state degrades stochastically until an expensive restorative action is taken, while all assets share a fixed intervention budget.
+For a large numbers of assets, deriving an optimal policy for this joint POMDP is computationally intractable.
+To tackle this challenge, we prove that the value function of the associated belief-MDP is budget-concave, which allows an efficient two-step approach to finding a near-optimal policy.
+First, we approximate the optimal cross-component budget split via a random-forest surrogate of each single-component value function.
+Second, we solve each resulting budget-constrained single-component POMDP with an oracle-guided meta-trained Proximal Policy Optimization (PPO) policy: value-iteration on the fully observable counterpart yields an oracle that shapes the PPO update and greatly accelerates learning.
+We validate our method through experiments in two disparate domains: (i) preventive maintenance for a large-scale building infrastructure containing 1,000 components, and (ii) portfolio risk management under debit-only loss-budget constraints, where each asset’s latent budget depletes with market losses and can only be replenished through costly recapitalization.
+Results show that our method consistently achieves longer component survival times and enhanced portfolio viability than both baseline heuristics and vanilla PPO.
+Furthermore, our approach maintains linear scalability in solution time with respect to the number of components.
 1 Introduction
-Partially Observable Markov Decision Processes (POMDPs) provide an efficient framework for modeling a wide range of real-world sequential decision-making problems (Cassandra 1998; Bravo, Leiras, and Cyrino Oliveira 2019).
-Numerous methods have been developed to compute optimal policies for POMDPs, including Monte-Carlo
+Partially Observable Markov Decision Processes (POMDPs) offer a principled framework for sequential decision making under uncertainty regarding the true state of the system  Cassandra (1998); Bravo et al.
+(2019).
+Solving POMDPs is computationally challenging, leading to the development of various solvers, including Monte-Carlo tree search Katt et al.
+(2017), reinforcement-learning variants Singh et al.
+(2021), and diverse approximation schemes Kearns et al.
+(1999).
+Many application domains share a monotonic structure, where the latent state of individual components degrades stochastically over time unless a costly restorative action is taken.
+Canonical examples include online advertising  Boutilier & Lu (2016), inventory replenishment Shin & Lee (2015), and sequential repair or maintenance scheduling for physical assets Miehling & Teneketzis (2020); Bhattacharya et al.
+(2021).
 Figure 1: Condition of infrastructure component over time.
@@
 (b) Violin-plot showing distribution of component condition for different time steps.
-methods (Katt, Oliehoek, and Amato 2017), reinforcement learning approaches (Singh et al.
-2021), and various approximation techniques (Kearns, Mansour, and Ng 1999).
-One notable application of POMDPs is in infrastructure maintenance or sequential repair problems, where the system state represents the condition of infrastructure components, which is only partially observable (Bhattacharya et al.
-2021).
-Over time, the state of these components stochastically deteriorates unless a restorative action, such as a repair, is performed.
-Figure 1 shows this stochastic decrease and also shows the probability distribution of the condition of a sample infrastructure component for different time steps.
-The state evolution of the POMDP, thus, exhibits monotonic properties (Miehling and Teneketzis 2020) and we define such POMDPs as monotonic POMDPs.
-Previously, Bhattacharya et al.
-(2020) proposed an optimal inspection and repair scheduling policy for a pipeline, which is a single-component POMDP.
-However, infrastructure systems naturally have multiple components (Daulat et al.
-2024).
+Figure 1 shows this stochastic decline and the probability distribution of a sample component’s condition at successive time steps.
+While prior work, such as Bhattacharya et al.
+(2020), has addressed optimal policies for single-component systems, real-world systems—from building portfolios to exchange-traded-fund (ETF) baskets—naturally involve many such components Daulat et al.
+(2024).
 In this paper, we address the challenge of computing approximately optimal policies for budget-constrained multi-component monotonic POMDPs.
-Each component POMDP operates independently in terms of transition probabilities, but they are collectively constrained by the shared budget.
-There has been substantial work on solving budget constrained POMDPs (Lee et al.
-2018; Undurti and How 2010; Khonji, Jasour, and Williams 2019).
-However, the complexity of these algorithms is often exponential in the number of states of the POMDP, and consequently would be exponential in the number of components for a multi-component POMDP.
-Thus, they become computationally intractable for multi-component POMDPs with a large number of components.
+We assume that each component POMDP operates independently in terms of transition probabilities, but they are collectively constrained by the shared budget.
+Substantial work has been done to solve budget-constrained POMDPs Lee et al.
+(2018); Undurti & How (2010); Khonji et al.
+(2019).
+However, the complexity of these algorithms is often exponential in the number of states of a single POMDP.
+For a multi-component POMDP, where the overall state space is the Cartesian product of individual component state spaces, this complexity consequently becomes exponential in the number of components.
+Thus, these methods become computationally intractable for multi-component POMDPs with a large number of components.
+A key challenge in solving budget-constrained multi-component POMDPs is how to optimally allocate the shared budget across the multiple components.
 In Vora et al.
 (2023), the authors propose a welfare-maximization method for solving budget-constrained multi-component POMDPs.
-However, the method in that paper requires generating optimal policies for multiple budget values for every component POMDP to get the optimal budget allocation, i.e., optimal distribution of the shared budget among the components POMDPs.
+However, the method in that paper requires generating optimal policies for multiple budget values for every component POMDP to get the optimal budget allocation.
 Hence, it cannot be scaled to a large number of components.
-To address this scalability issue, we propose a computationally efficient meta-reinforcement learning approach for solving budget-constrained multi-component monotonic POMDPs.
-First, we approximate the optimal distribution of budget among the individual component POMDPs using a random forest budget allocation strategy, thereby decoupling the problem into independent components.
-The random forest model is trained to approximate the value function of a component POMDP as a function of the budget, leveraging the concavity of the value function to frame the budget allocation as a concave maximization problem, which guarantees a computationally feasible solution.
-Subsequently, we apply the oracle policy-guided meta-trained Proximal Policy Optimization (PPO) algorithm, trained across various components and budget values, to compute the approximately optimal policy for each component-budget pair.
-The oracle policy for the meta-PPO is derived from value iteration applied to the corresponding component MDP.
-To validate the efficacy of our proposed approach, we examine a maintenance scenario involving an administrative building with 1000 infrastructure components and a fixed maintenance budget.
-We compare the performance of our approach, in terms of the survival time of the components, against the oracle policy, a currently-used baseline policy, and a vanilla meta-PPO algorithm.
-We also perform a computational complexity analysis of our algorithm for varying number of components and show that the proposed approach is linear in the number of components.
-In summary, the key contributions of this paper are:
+Our insight.
+The primary computational bottleneck in solving budget-constrained multi-component POMDPs is the coupling induced by the shared budget.
+If that budget could be split a-priori, the joint POMDP would factor into $n$ independent, single-component problems solvable in parallel.
+To enable this decomposition, we prove that the optimal value function of a single monotonic POMDP is concave in its allocated budget.
+This budget-concavity lets us decouple first, optimise second:
+  1. (1)
+    Budget allocation. We maximize a concave surrogate of the value function, estimated with a random-forest regressor, to distribute the global budget across components; and
+  2. (2)
+    Component policies.
+With budgets fixed, we learn a near-optimal policy for each component–budget pair using an oracle-guided, meta-trained Proximal Policy Optimization (PPO) agent, where the oracle is obtained by value iteration on the fully observable counterpart.
+The result is a scalable solution whose runtime grows linearly with the number of components while retaining strong performance guarantees.
+Contributions.
   1. 1.
-    We develop an oracle policy-guided meta-PPO algorithm for solving budget-constrained single-component monotonic POMDPs.
+    Theory.
+We prove budget-concavity of the optimal value function for monotonic POMDPs.
+While prior works implicitly assume and use this budget-concavity, our work provides the first general structural guarantee that formally links budget availability to expected return.
   2. 2.
-    We introduce a random forest budget allocation strategy for optimally distributing the budget among the component POMDPs of a budget-constrained multi-component POMDP.
+    Algorithms.
+We introduce (i) a random-forest budget-allocation module that exploits concavity for fast global optimization, and (ii) an oracle-guided meta-PPO solver for each single-component POMDP.
   3. 3.
-    We demonstrate the efficacy of our approach through a large-scale real-world infrastructure maintenance scenario with 1000 infrastructure components.
+    Empirical evidence.
+On two domains—preventive maintenance of a 1000-component building and portfolio loss-budget management with recapitalization—we outperform baseline heuristics and vanilla PPO, whilst solution time of the proposed approach scales linearly in the number of components.
   4. 4.
-    We show the scalability of our algorithm through a computational complexity analysis for varying number of components.
+    Complexity analysis. We provide a detailed runtime study confirming linear growth in wall-clock time as components increase from $n=10$ to $n=1000$.
+The remainder of the paper is organized as follows.
+Section 2 surveys related work on budget-constrained POMDPs and large-scale maintenance or portfolio problems.
+Section 3 formalises the budget-constrained multi-component monotonic POMDP.
+Section 4 presents our solution pipeline: (i) Subsection 4.1 proves budget–concavity of the single-component value function; (ii) Subsection 4.2 exploits this structure to allocate the global budget via a random-forest surrogate; and (iii) Subsection 4.3 derives an oracle-guided meta-PPO policy for each component and composes them into the overall controller.
+Section 5 reports empirical results on infrastructure maintenance and financial loss-budget management, and Section 6 concludes with key findings and future directions.
 2 Preliminaries and Related Work
 2.1 Partially Observable Markov Decision Processes
-A discrete-time finite-horizon Partially Observable Markov Decision Process (POMDP) (Cassandra, Kaelbling, and Littman 1994) $M$ is defined by the 7-tuple $(\mathcal{S},A,T,\Omega,O,R,H)$, which denotes the state space, action space, state transition function, observation space, observation function, reward function and planning horizon, respectively.
+A discrete-time finite-horizon Partially Observable Markov Decision Process (POMDP) Cassandra et al.
+(1994) $M$ is defined by the 7-tuple $(\mathcal{S},A,T,\Omega,O,R,H)$, which denotes the state space, action space, state transition function, observation space, observation function, reward function and planning horizon, respectively.
 In a POMDP, the agent does not have direct access to the true state of the environment.
 Instead, the agent may maintain a belief state, representing a probability distribution over $\mathcal{S}$.
-This belief is updated based on the received observation using Bayes’ rule (Araya et al.
-2010).
+This belief is updated based on the received observation using Bayes’ rule Araya et al.
+(2010).
 2.2 POMDP Solution Methods
-Computing optimal policies for a POMDP is generally PSPACE-complete (Mundhenk et al.
-2000; Vlassis, Littman, and Barber 2012).
-Thus, to address the computational intractability of solving POMDPs, various approximation methods have been widely used (Poupart and Boutilier 2002; Pineau et al.
-2003; Roy, Gordon, and Thrun 2005).
-Several reinforcement learning approaches have also been developed for computing approximate POMDP solutions (Azizzadenesheli, Lazaric, and Anandkumar 2016; Igl et al.
-2018).
-However these methods become computationally intractable when faced with the high dimensionality and shared resource constraints of budget-constrained multi-component monotonic POMDPs such as those considered in this paper.
+Computing optimal policies for a POMDP is generally PSPACE-complete Mundhenk et al.
+(2000); Vlassis et al.
+(2012).
+Thus, to address the computational intractability of solving POMDPs, various approximation methods have been widely used Poupart & Boutilier (2002); Pineau et al.
+(2003); Roy et al.
+(2005).
+Several reinforcement learning approaches have also been developed for computing approximate POMDP solutions Azizzadenesheli et al.
+(2016); Igl et al.
+(2018).
+However, these methods become computationally intractable when faced with the high dimensionality and shared resource constraints of budget-constrained multi-component monotonic POMDPs such as those considered in this paper.
 2.3 Consumption MDPs and Budgeted POMDPs
-The integration of budget or resource constraints into Markov Decision Processes (MDPs) has been previously studied under the frameworks of Consumption MDPs (Blahoudek et al.
-2020) and Budgeted POMDPs (Vora et al.
-2023).
+The integration of budget or resource constraints into Markov Decision Processes (MDPs) has been previously studied under the frameworks of Consumption MDPs Blahoudek et al.
+(2020) and Budgeted POMDPs Vora et al.
+(2023).
 However, the algorithm proposed in Blahoudek et al.
@@
 The transition probability function for each component $i$ is:
-$T^{i}(s^{i}_{k},a^{i}_{k},s^{i}_{k+1})=\begin{cases}p_{1}^{i}(s^{i}_{k},a^{i}_{k},s^{i}_{k+1}),&\text{if }a^{i}_{k}=m^{i}\text{ and }\\ &s^{i}_{k}\leq s^{i}_{k+1}\leq\bar{s},\\ p_{2}^{i}(s^{i}_{k},a^{i}_{k},s^{i}_{k+1}),&\text{if }a^{i}_{k}\in\{d^{i},q^{i}\}\\ &\text{and }s^{i}_{k+1}\leq s^{i}_{k},\\ 1,&\text{if }a^{i}_{k}\in A^{i}\text{ and }\\ &s^{i}_{k+1}=0=s^{i}_{k},\\ 0,&\text{otherwise}.\end{cases}$
+[…]
 Here, action $m^{i}$ is a restorative action that increases the state value, with the increase being upper bounded by $\bar{s}$.
@@
 The primary objective of this paper is to determine a policy $\pi^{*}$ for this multi-component monotonic POMDP over a horizon $H$, that maximizes the sum of expectations of the individual times before reaching the absorbing state for each component, while adhering to the total budget $B$.
-We denote this maximal time $k$ by $T_{max}=\sum_{i=1}^{n}T^{i}_{max}$, where $T^{i}_{max}$ denotes the corresponding maximal time for component $i$.
+We denote this maximal time $k$ by $T_{max}=\sum_{i=1}^{n}T^{i}_{max}$, where $T^{i}_{max}$ denotes the corresponding maximal time for component ii.
 Mathematically, the problem can be formulated as:
 $\max_{\pi}\left(\sum_{i=1}^{n}\mathbb{E}[T^{i}_{max}(\pi)]\right)$ (2)
-[…]
+$\text{s.t.}\sum_{k=0}^{H}c_{a_{k}}(\pi)\leq B.$
 In this formulation, $\pi$ represents the policy, and both $T^{i}_{max}$ and $c_{a_{k}}$ depend on $\pi$.
@@
 Figure 2 presents an architectural overview of our proposed approach.
-We first input information about all $n$ component POMDPs into a pre-trained random forest regressor to get the $T^{i}_{max}$ for each component $i$ as a function of the budget allocated to the component.
-Next, we propose an appropriate allocation of the shared budget $B$ among the component POMDPs by solving a constrained maximization problem.
-We then compute the oracle policy for each component POMDP and budget pair through value iteration applied to the corresponding component MDP.
-Finally, using these oracle policies and a meta-trained reinforcement learning (RL) agent, we solve each component POMDP with respect to the allocated budget and consequently propose a policy for the multi-component POMDP.
+The key idea is to decouple first, optimize second.
+Allocating the shared budget as a first step of planning shrinks the original large joint POMDP into $n$ independent single-component POMDPs.
+Each of these single-component POMDPs now operates with its own fixed budget cap, which is determined by the initial allocation.
+This transformation converts a problem that is intractable for $n\!\gg\!1$ into $n$ modest ones that can be solved in parallel.
+We organize the section accordingly:
+  • Structural Result: Budget Concavity (Section 4.1): We prove that each component’s value function is concave in its budget.
+  • Stage 1: Budget Allocation (Section 4.2): Leveraging concavity, we fit a random-forest surrogate of the value function and solve a tractable concave maximization problem to distribute the shared global budget across components.
+  • Stage 2: Oracle-Guided Meta-PPO (Section 4.3): With budgets fixed, we learn near-optimal policies for each component–budget pair (with respect to that component’s allocated budget and local POMDP) using an oracle-guided, meta-trained PPO agent, then compose these into the overall multi-component policy.
 Note that an alternate allocation strategy could involve redistributing the budget at every time step during planning.
-However, such a method would be computationally more expensive, due to the repeated computation of the allocation, as compared to our proposed a priori budget distribution approach.
-4.1 Oracle-Guided RL for Budget-Constrained Monotonic POMDPs
-To derive the optimal policy for a budget-constrained single-component monotonic POMDP, we propose an oracle policy-guided reinforcement learning algorithm.
-Adherence to the budget constraints is achieved using the budget-constrained POMDP (bPOMDP) framework proposed in Vora et al.
-(2023).
-In a bPOMDP, the state at each time step includes an additional fully observable component representing the cost incurred up to that point.
+However, such a method would be computationally more expensive than our proposed approach due to the repeated computation of the allocation.
+4.1 Budget–Concavity of the Value Function
+We first show that the optimal value function of a single-component belief-MDP—derived from a monotonic POMDP with random, nonnegative action costs—is concave in the budget variable $B$ for any finite planning horizon $H\geq 0$.
+Setting
+We consider a POMDP defined by the tuple $\langle S,A,T,R,\Omega,O,\gamma,\mathcal{C}\rangle$, where:
+  • $S$ is a finite state space and AA is a finite action space, as defined in Section 3.
+  • $T(s^{\prime}|s,a)$ is the transition kernel (1); $R(s,a)$ is the reward function.
+  • $\Omega$ is the observation space; $O(o|s^{\prime},a)$ is the observation model as defined in Section 3.
+  • $\gamma\in[0,1)$ is the discount factor.
+  • $\mathcal{C}(s,a)$ is the distribution of a random, nonnegative cost incurred by taking action $a$ in state $s$. The specific cost, $c$, is drawn from this distribution.
+Following standard practice in POMDP literature Cassandra (1998), we reformulate this POMDP as a belief-MDP with state $(b,B)$, where $b\in\Delta(S)$ is the belief (posterior distribution over hidden states), and $B\geq 0$ is a remaining budget.
+The reward at belief $b$ under action $a$ is:
+$\rho(b,a):=\sum_{s\in S}b(s)R(s,a),$
+and the budget evolves as $B\mapsto B-c$, where $c$ is a realization from the random cost $C(s,a)$ under the current belief.
+To prove the budget-concavity of the value function for this belief-MDP, we first establish two foundational properties concerning concavity.
+These lemmas demonstrate how concavity is preserved under common mathematical operations relevant to dynamic programming.
+Lemma 1 (Concavity under Affine Shift).
+If $f(x)$ is concave on an interval II, then f(xl)f(x-l) is concave on the interval $\{y\mid y=x+l,x\in I\}$ for any constant $l$.
+Proof.
+This lemma is a standard result in convex analysis Boyd & Vandenberghe (2004). ∎
+Lemma 2 (Expectation Preserves Concavity).
+Let $f(B,\xi)$ be concave in $B$ for every realization $\xi$. If $\xi$ is a random variable following an arbitrary probability distribution, then $\mathbb{E}_{\xi}[f(B,\xi)]$ is also concave in $B$.
+Proof.
+Fix any $B_{1},B_{2}\in\mathbb{R}$ and any $\lambda\in[0,1]$. Let $g(B)=\mathbb{E}_{\xi}[f(B,\xi)]$. Then
+$g(\lambda B_{1}+(1-\lambda)B_{2})=\mathbb{E}_{\xi}\left[f\big{(}\lambda B_{1}+(1-\lambda)B_{2},\xi\big{)}\right].$
+Since $f(B,\xi)$ is concave in $B$ for each $\xi$, we have
+$f\big{(}\lambda B_{1}+(1-\lambda)B_{2},\xi\big{)}\geq\lambda f(B_{1},\xi)+(1-\lambda)f(B_{2},\xi)$
+for all $\xi$. Taking expectations on both sides yields
+$\mathbb{E}_{\xi}\left[f(\lambda B_{1}+(1-\lambda)B_{2},\xi)\right]\geq\lambda\mathbb{E}_{\xi}\left[f(B_{1},\xi)\right]+(1-\lambda)\mathbb{E}_{\xi}\left[f(B_{2},\xi)\right],$
+that is,
+$g(\lambda B_{1}+(1-\lambda)B_{2})\geq\lambda g(B_{1})+(1-\lambda)g(B_{2}).$
+Thus, $g$ is concave in $B$. ∎
+Having established these fundamental properties regarding the preservation of concavity under affine shifts and expectations, we will now prove that the optimal value function of a monotonic POMDP is concave with respect to the available budget.
+Theorem 3 (Budget Concavity).
+For any fixed belief $b\in\Delta(S)$ and horizon $H\geq 0$, the value function $V_{H}(b,B)$ is concave in $B$ on $0,\infty)$.
+Proof.
+We proceed by mathematical induction on $H$.
+Base Case ($H=0$). At horizon zero, there are no rewards:
+$V_{0}(b,B)=0\quad\text{for all }b\in\Delta(S),\ B\geq 0.$
+Function $V_{0}$ is thus trivially concave.
+Inductive Hypothesis. Suppose that for some $H\geq 0$, the function $V_{H}(b,B)$ is concave in $B$ for every belief $b$.
+Inductive Step. We aim to prove that $V_{H+1}(b,B)$ is concave in $B$ for all $b$. The Bellman equation in the belief-MDP is:
+$V_{H+1}(b,B)=\max_{a\in A}\left\{\rho(b,a)+\gamma\,\mathbb{E}_{o,c\mid b,a}\left[V_{H}(b^{\prime},B-c)\right]\right\},$
+where $b^{\prime}$ is the updated belief after taking action $a$ and observing $o$.
+The expectation $\mathbb{E}_{o,c\mid b,a}$ is taken over the random observation $o$ and cost $c$ given the current belief $b$ and chosen action $a$.
+Define the inner expectation as:
+$g(a,b,B):=\mathbb{E}_{o,c\mid b,a}\left[V_{H}(b^{\prime},B-c)\right].$
+Apply Lemma 1 and inductive hypothesis to assert that $B\mapsto V_{H}(b^{\prime},B-c)$ is concave for each $(o,c)$.
+Then Lemma 2 implies that $g(B)$, being the expectation over such functions, is also concave.
+Therefore, the $Q$-value
+$Q_{H+1}(b,B,a):=\rho(b,a)+\gamma g(B)$
+is concave in $B$ for each $a$.
+Finally, the value function is
+$V_{H+1}(b,B)=\max_{a\in A}Q_{H+1}(b,B,a),$
+which is the pointwise maximum of finitely many concave functions, and hence concave itself. ∎
+4.1.1 Relating $\mathbb{E}[T_{\max}]$ to the Value Function
+We proved the budget-concavity of the value function $V_{H}(s,B)$ in Theorem 3.
+In our problem setting (2), however, we aim to maximize the expected time to failure $E[T_{\max}]$.
+Specifically, in many practical applications such as preventive maintenance or portfolio management, the objective can be naturally framed as maximizing the expected time until a critical failure occurs or a budget is exhausted.
+We now show how the concavity property extends to $E[T_{\max}]$, which serves as the objective function for our initial budget allocation stage.
+Lemma 4 (Expected-time equivalence).
+Consider the reward function
+$R(s,a)=\begin{cases}1,&s\neq 0,\\ 0,&s=0,\end{cases}$ (4)
+. Let $V(s,B)$ be the corresponding optimal value function. Denote by $\mathbb{E}[T_{\max}(B)]$ the expected time to reach the absorbing state $s=0$ under the optimal budget-feasible policy. Then
+$V(s,B)=\mathbb{E}[T_{\max}(B)].$
+Proof.
+Under reward scheme (4) each non-absorbing step contributes exactly 11 to the return; steps in state 0 contribute 0. Hence, for any budget-feasible policy $\pi$,
+$\text{Total reward}=\mathbb{E}\Bigl{[}\sum_{t=0}^{H}\mathbf{1}\{s_{t}\neq 0\}\Bigr{]}=\mathbb{E}[T_{\max}(\pi)].$
+Taking the maximum over all budget-feasible policies yields $V(s,B)=\mathbb{E}[T_{\max}(B)]$. ∎
+Corollary 5.
+Because $B\mapsto V(s,B)$ is concave by Theorem 3, the expected absorption time $\mathbb{E}[T_{\max}(B)]$ is also concave in the allocated budget $B$.
+4.2 Random Forest Approach for Optimal Budget Allocation
+By Corollary 5, the expected maximal survival time $\mathbb{E}[T_{\max}(B)]$ is a concave function of the budget allocated to a single component.
+This structural property lets us treat budget splitting across $n$ components as a concave maximization problem—one that is both tractable and amenable to surrogate modeling.
+Each component evolves independently but competes for the shared budget, rendering the components weakly coupled.
+While reinforcement learning algorithms have made significant advances, they often face challenges when scaling to the extremely large state and action spaces characteristic of multi-component systems Sutton & Barto (2018).
+To address this scalability issue, our remedy is an a-priori budget distribution that decouples the system.
+Concretely, for component $i$ we approximate the concave map $B\mapsto\mathbb{E}[T_{\max}^{i}(B)]$ by the exponential surrogate
+$\widetilde{T}^{i}_{\max}(B)\;=\;\alpha^{i}\,e^{\beta^{i}B}+\gamma^{i},$ (5)
+where $(\alpha^{i},\beta^{i},\gamma^{i})$ are constants.
+While many other concave functions could be used to model $\tilde{T}^{i}_{\max}$, we empirically observe that the exponential function provides a good fit for the data (see Appendix A).
+We use a random forest regressor Breiman (2001) to estimate the parameters of this exponential function.
+The training dataset for this model is obtained via non-linear least squares regression on multiple $(\mathbb{E}[T_{\max}],b)$ pairs for various budget-constrained single-component monotonic POMDPs.
+The input to this model includes specific statistics related to the POMDP’s transition function, which are the expected time to reach state 0 without repairs, $\mathbb{E}[T]$, and the variance of this expected time, $\sigma^{2}_{\mathbb{E}[T]}$, as well as the various actions costs.
+Let $b^{i}$ denote the budget assigned to component $i$ and $\widetilde{T}^{i}_{\max}$ its surrogate survival time. The allocation problem becomes
+$\max_{b^{1:n}}$ $\sum_{i=1}^{n}\widetilde{T}^{i}_{\max}(b^{i})$ (6)
+s.t. $\sum_{i=1}^{n}b^{i}\;\leq\;B,\qquad b^{i}\;\geq 0\;\;\forall i,$
+a concave maximization with linear constraints.
+Because each surrogate in (5) is concave, the problem is globally tractable and we solve it with off-the-shelf convex optimizers.
+Solving (6) yields the approximately optimal budget allocation among the individual components.
+The next subsection shows how an oracle-guided meta-PPO agent learns the individual component policies given this budget allocation.
+4.3 Oracle-Guided RL for a Budget-Constrained Single Component
+Given the per-component budgets $b^{i}$ obtained in Section 4.2, we now derive a near-optimal control policy for each single-component budget-constrained monotonic POMDP.
+We adopt the budget-augmented POMDP (bPOMDP) formalism of Vora et al.
+(2023), in which the state includes an additional, fully-observable coordinate that tracks cumulative cost.
 The oracle policy is denoted as $\pi_{\text{oracle}}$ and is obtained by solving the corresponding MDP using value iteration.
 For a single-component monotonic POMDP with budget $B$, the corresponding MDP has an action space $\mathcal{A}_{\text{MDP}}=\{d,m\}$, identical transition probabilities as the POMDP, and full observability of the state.
-We use the Proximal Policy Optimization (PPO) algorithm (Schulman et al.
-2017) in conjunction with the oracle policy to solve the single-component budget-constrained POMDP.
-Doing so allows us to treat the problem as that of determining the optimal inspection policy.
-At each time step, the PPO agent decides whether to perform inspection or not.
-If the agent decides to take the inspection action, it receives accurate information about the true state of the POMDP.
-On the other hand, if the agent decides not to inspect, then the action is chosen according to the oracle policy.
-Hence, the action space for the PPO agent is reduced to $\{q,\neg q\}$.
+We then train a Proximal Policy optimization (PPO) agent Schulman et al.
+(2017) that queries this oracle selectively: at each time step it chooses either to inspect ($q$) or to defer ($\neg q$), in which case the action recommended by the oracle is executed.
 Since the full state is not observable in a POMDP, we utilize the belief $b_{s}$ for planning.
@@
 For our work, we empirically observe that using the expected belief $\bar{b_{s}}$ and the variance of the belief $\sigma^{2}_{b_{s}}$ suffices for planning.
-Hence, for the proposed oracle policy-guided PPO agent, the state at time instant $k$ is given by the vector $[\bar{b}_{s_{k}},c_{k},\sigma^{2}_{b_{s_{k}}}]$.
+Hence, for the proposed oracle policy-guided PPO agent, the state at time instant $k$ is given by the vector $\bar{b}_{s_{k}},c_{k},\sigma^{2}_{b_{s_{k}}}]$.
 Furthermore, the reward function is defined as follows:
 $R(s_{k},c_{k},a_{k})=\begin{cases}r_{1}<0,&\text{if }c_{k}>B,\\ r_{2}<0,&\text{if }\left\lfloor{\bar{b}_{s_{k}}}\right\rfloor=0,\\ r_{3}=\frac{k}{H}-\alpha|\bar{b}_{s_{k}}-s_{k}|,&\text{if }\bar{b}_{s_{k}},c_{k}>0,\end{cases}$
-[…]
-4.2 Random Forest Approach for Optimal Budget Allocation
-In the previous subsection, we detailed the methodology for deriving the optimal policy for a single-component budget-constrained monotonic POMDP.
-We now extend this to an $n$-component POMDP described in Section 3.
-Each component in this multi-component POMDP operates independently in terms of transition probabilities, but they share a common total budget, rendering them weakly coupled.
-While reinforcement learning algorithms have made significant advances, they often face challenges when scaling to the extremely large state and action spaces characteristic of multi-component systems (Sutton and Barto 2018).
-In the case of a budget-constrained $n$-component POMDP, directly applying a single-component solution approach would result in an exponentially large state space and action space, making such an approach impractical for large-scale problems.
-To address this scalability issue, we propose an a priori budget distribution strategy to decouple the components, thereby reducing the size of the state and action spaces.
-Under certain assumptions, the expected maximal time to reach the absorbing state, $\mathbb{E}[T_{\max}]$, for a budget-constrained single-component monotonic POMDP, is concave with respect to the allocated budget $b$ (Vora et al.
-2023).
-For a given component $i$ and allocated budget $b^{i}$, let $\tilde{T}^{i}_{\max}$ be the approximate expected maximal time as a function of the budget, modeled using an exponential function:
-$\tilde{T}^{i}_{\max}=\alpha^{i}e^{\beta^{i}b^{i}}+\gamma^{i},$ (3)
-where $\alpha^{i},\beta^{i},\gamma^{i}\in\mathbb{R}$ are constants.
-While many other concave functions could be used to model $\tilde{T}^{i}_{\max}$, we empirically observe that the exponential function provides a good fit for the data.
-We use a random forest regressor (Breiman 2001) to estimate the parameters of this exponential function.
-The training dataset for this model is obtained via non-linear least squares regression on multiple $(\mathbb{E}[T_{\max}],b)$ pairs for various budget-constrained single-component monotonic POMDPs.
-The input to this model includes specific statistics related to the POMDP’s transition function, which are the expected time to reach state 0 without repairs, $\mathbb{E}[T]$, and the variance of this expected time, $\sigma^{2}_{\mathbb{E}[T]}$, as well as the cost of the various actions.
-For an $n$-component POMDP with total budget $B$, let the budget allocated to the $i$-th component be $b^{i}$, and the approximate expected maximal time to reach the absorbing state for this budget be $\tilde{T}^{i}_{\max}$.
-Figure 3: Performance comparison of oracle policy, oracle-guided meta-PPO, realistic baseline and vanilla meta-PPO.
-(a) Comparison of $\hat{T}_{max}$ values for all 1000 components across different budget values allocated to each component.
-(b) Comparison of average number of repairs performed by the agent under each of the four policies.
-(c) Comparison of average total cost incurred by the agent over the planning horizon for each of the four policies.
-We then formulate the following constrained maximization problem:
-$\max_{b^{i}}\sum_{i=1}^{n}\tilde{T}^{i}_{\max}$ (4)
-[…]
-Solving (4) yields the approximately optimal budget allocation among the individual components.
-Since we assume $\tilde{T}^{i}_{\max}$ to be concave in $b^{i}$, the optimization problem becomes a concave maximization problem and is thus guaranteed to have a computationally tractable solution.
+where $|r_{1}|>|r_{2}|>|r_{3}|$ for all $k$, $0<\alpha<1$ and $\left\lfloor{.}\right\rfloor$ denotes the floor function.
+This reward function imposes substantial negative rewards for exceeding the budget $B$ and allowing the state $s_{k}$ to reach 0.
+Additionally, at each time step, the agent receives a positive reward proportional to the time step for maintaining $s_{k}$ above zero and incurs a penalty proportional to the absolute error between the expected belief and the true state.
+As a result, the agent gets higher rewards for keeping $s_{k}>0$ for a longer time and is heavily penalized when the expected belief deviates significantly from the true state.
+It is crucial to note that during training, the agent relies solely on the observed reward signals, without access to the true state.
 4.4 Optimal Policy for Multi-Component Monotonic POMDPs
-We now integrate the approaches described in the previous subsections to compute the optimal policy for an $n$-component POMDP, where $n$ is substantially large.
-Utilizing the random forest regressor from Section 4.2, we efficiently approximate $\mathbb{E}[T_{\max}]$ for each component $i$.
-Additionally, we meta-train the oracle-guided PPO agent by constantly updating the policy network’s parameters over a randomly selectied a subset of components and budget values.
+We now integrate the approaches described in Section 4.2 and Section 4.3 to compute the optimal policy for an $n$-component POMDP, where $n$ is substantially large.
+Utilizing the random forest regressor, we efficiently approximate $\mathbb{E}[T_{\max}]$ for each component ii.
+Additionally, we meta-train the oracle-guided PPO agent by continuously updating the policy network’s parameters over a randomly selected subset of components and budget values.
 This approach allows the agent to generalize across components.
@@
 5 Implementation and Evaluation
+In this section, we empirically validate our proposed framework on two disparate domains.
+The first domain, which we call the infrastructure scenario, involves preventive maintenance for a large-scale building comprising 1000 independent components whose latent condition stochastically degrades over time; our goal is to allocate a finite maintenance budget to maximize the expected survival time of all components.
+The second domain, the financial loss-budget scenario, addresses portfolio risk management using daily price data for S&P 500 constituents, where each asset is endowed with a debit-only loss budget that depletes under negative returns and can be replenished only through costly recapitalization.
+In the infrastructure scenario, we compare our oracle-guided meta-PPO approach against baseline heuristics, vanilla PPO, and an idealized oracle policy, reporting results on survival time, cost efficiency, and computational scalability across a range of budget levels.
+In the financial loss-budget scenario, we focus on analyzing the learned recapitalization policy and assessing the generalizability and window robustness of proposed oracle-guided meta-PPO.
+5.1 Implementation and Evaluation for Infrastructure Scenario
 In this section, we evaluate the efficacy of the proposed methodology for determining the optimal policy for a very large multi-component budget-constrained POMDP.
@@
 We consider an administrative building comprising 1000 infrastructure components, including roofing elements, water fountains, lighting systems, and boilers.
-Each component’s health is quantified by the Condition Index (CI) (Grussing, Uzarski, and Marrano 2006), which ranges from 0 to 100.
-For each infrastructure component, we utilize historical CI data to generate the transition probabilities for the corresponding POMDP, modeled using the Weibull distribution (Grussing, Uzarski, and Marrano 2006).
+Each component’s health is quantified by the Condition Index (CI) Grussing et al.
+(2006), which ranges from 0 to 100.
+For each infrastructure component, we utilize historical CI data to generate the transition probabilities for the corresponding POMDP, modeled using the Weibull distribution Grussing et al.
+(2006).
 We use the weibull_min class from the scipy.stats module in Python to simulate the CI transitions over time.
@@
 This POMDP has roughly $10^{2000}$ states and $3^{1000}$ actions.
-5.1 Analysis of Maintenance Policy
+Figure 3: Performance comparison of oracle policy, oracle-guided meta-PPO, realistic baseline and vanilla meta-PPO.
+(a) Comparison of $\hat{T}_{max}$ values for all 1000 components across different budget values allocated to each component.
+(b) Comparison of average number of repairs performed by the agent under each of the four policies.
+(c) Comparison of average total cost incurred by the agent over the planning horizon for each of the four policies.
+5.1.1 Analysis of Maintenance Policy
 We begin by analyzing the performance of the maintenance policy derived using the proposed oracle-guided meta-PPO strategy for a single-component POMDP representing a component $i$ of the 1000 components.
@@
   1. 1.
-    A heuristic policy often used in practice (Lam and Yeh 1994; Straub 2004) where the agent performs inspections at regular intervals and repairs the component when its expected belief about the Condition Index (CI) falls below a predefined threshold.
+    A heuristic policy often used in practice Lam & Yeh (1994); Straub (2004) where the agent performs inspections at regular intervals and repairs the component when its expected belief about the Condition Index (CI) falls below a predefined threshold.
 We chose an inspection interval of 5 steps and a repair threshold of 15 after extensive experiments with intervals ranging from 1 to 10 steps and repair thresholds from 5 to 50.
@@
 These results demonstrate the value of incorporating an oracle policy into the training of a reinforcement learning agent.
-5.2 Analysis of Budget Allocation
+Figure 4: Performance of random forest model for predicting the value of parameter $\beta$ for a test dataset of 200 components.
+The horizontal axis represents parameter values obtained via non-linear least squares and vertical axis represents predicted values.
+The dotted line represents the $y=x$ line, i.e., perfect predictions.
+5.1.2 Analysis of Budget Allocation
 Next, we demonstrate the effectiveness of our random forest-based budget allocation strategy.
 We compare it with a baseline approach that allocates budgets proportional to the ratio of a component’s replacement cost to its $\mathbb{E}[T]$.
-For a component $i$, we model its $\mathbb{E}[T^{i}_{\max}]$ using $\tilde{T}^{i}_{\max}$ as given in (3).
+For a component $i$, we model its $\mathbb{E}[T^{i}_{\max}]$ using $\tilde{T}^{i}_{\max}$ as given in (5) (see Appendix A for justification of this exponential form).
 The parameters $\alpha^{i}$ and $\gamma^{i}$ can be estimated directly by considering the boundary conditions: $\gamma^{i}$ is estimated by substituting $b^{i}=0$, representing the scenario where no budget is available, and $\alpha^{i}$ is determined by substituting $b^{i}=\infty$, corresponding to the scenario of unlimited budget, where the supremum of $T^{i}_{\max}$ ($\sup_{b^{i}}T^{i}_{\max}=H=100$) is reached.
@@
 The training dataset is created by performing non-linear least squares regression on 11 distinct ($T^{i}_{max},b^{i}$) pairs each for 800 components.
-These pairs correspond to the run-averaged $T^{i}_{\max}$ values and the respective budget values $b^{i}$ from Section 5.1.
-The input to the random forest model is a vector consisting of the shape and scale factors of the Weibull distribution, which represent $\mathbb{E}[T]$ and $\sigma^{2}_{\mathbb{E}[T]}$, along with the replacement and inspection costs for a given component $i$.
-If a different distribution were used to model the transition probability, we would similarly extract the parameters, $\mathbb{E}[T]$ and $\sigma^{2}_{\mathbb{E}[T]}$, for inclusion in the input vector.
-Figure 4: Performance of random forest model for predicting the value of parameter $\beta$ for a test dataset of 200 components.
-The horizontal axis represents parameter values obtained via non-linear least squares and vertical axis represents predicted values.
-The dotted line represents the $y=x$ line, i.e., perfect predictions.
+These pairs correspond to the run-averaged $T^{i}_{\max}$ values and the respective budget values $b^{i}$ from Section 5.1.1.
+The input to the random forest model is a vector consisting of the shape and scale factors of the Weibull distribution, which represent $\mathbb{E}[T]$ and $\sigma^{2}_{\mathbb{E}[T]}$, along with the replacement and inspection costs for a given component ii.
+If a different distribution was used to model the transition probability, we would similarly extract the parameters, $\mathbb{E}[T]$ and $\sigma^{2}_{\mathbb{E}[T]}$, for inclusion in the input vector.
 Figure 4 shows the prediction performance of the random forest model for a test dataset of 200 components which were not encountered during training.
@@
 Table 1: $T_{max}$ (in steps), averaged over 100 runs, achieved under random forest and baseline budget allocations.
-Figure 5: Performance comparison of random forest-based budget allocation and baseline budget allocation for all 1000 components for an overall budget of 500,000 units.
 We quantify the performance of the random forest budget allocation and the baseline budget allocation algorithms by calculating the $T_{max}=\sum_{i}T^{i}_{max}$ and averaging it over 100 runs.
 For a fair comparison, these values are obtained using the oracle-guided meta-PPO approach for both allocation schemes.
+Table 1 shows the $T_{max}$ values achieved by both allocation approaches.
+The random forest budget allocation vastly outperforms the baseline approach.
+Furthermore, Figure 5 presents violin plots showing the distribution of the $T^{i}_{max}$ values achieved under the proposed and baseline budget allocations for all 1000 components.
+We observe that there are more components with higher $T^{i}_{max}$ values for the random forest budget allocation approach.
+Preliminary experiments on alternative objective formulations, such as the maxmin approach given by (3), also indicate that the proposed method consistently outperforms the baseline.
+Figure 5: Performance comparison of random forest-based budget allocation and baseline budget allocation for all 1000 components for an overall budget of 500,000 units.
+5.1.3 Analysis of Time Complexity
+Finally, we analyze the time complexity of our proposed approach for varying number of components NN. As mentioned earlier, our method comprises of four major steps:
+  1. 1.
+    Random Forest regression for estimating $\tilde{T}^{i}_{max}$ for each component ii.
+  2. 2.
+    Budget Allocation among components via constrained optimization.
+  3. 3.
+    MDP Value Iteration for each component-budget pair to obtain the corresponding oracle policy.
+  4. 4.
+    Oracle-Guided Meta-PPO to approximately solve each component POMDP.
 Number of Components Random Forest Budget Split Value Iteration Meta-PPO
@@
 Table 2: Time taken (in seconds) for running each process with varying numbers of components, averaged over 10 runs.
-Table 1 shows the $T_{max}$ values achieved by both allocation approaches.
-The random forest budget allocation vastly outperforms the baseline approach.
-Furthermore, Figure 5 presents violin plots showing the distribution of the $T^{i}_{max}$ values achieved under the proposed and baseline budget allocations for all 1000 components.
-We observe that there are more components with higher $T^{i}_{max}$ values for the random forest budget allocation approach.
-Preliminary experiments on alternative objective formulations, such as the maxmin approach given by (2), also indicate that the proposed method consistently outperforms the baseline.
-5.3 Analysis of Time Complexity
-Finally, we analyze the time complexity of our proposed approach for varying number of components $N$. As mentioned earlier, our method comprises of four major steps:
-  1. 1.
-    Random Forest regression for estimating $\tilde{T}^{i}_{max}$ for each component $i$.
-  2. 2.
-    Budget Allocation among components via constrained optimization.
-  3. 3.
-    MDP Value Iteration for each component-budget pair to obtain the corresponding oracle policy.
-  4. 4.
-    Oracle-Guided Meta-PPO to approximately solve each component POMDP.
+Figure 6: Log-log plot of computational complexity of the proposed approach for varying numbers of components.
 Table 2 presents the times taken for running each of the four processes, with different number of components.
@@
 This expectation is confirmed by the log-log plot of computational complexity shown in Figure 6.
-Our algorithm’s performance is thus significantly faster as compared to existing POMDP solvers which would be exponential in the number of states and thus doubly exponential in the number of components (Silver and Veness 2010), (Pineau et al.
-2003).
+Our algorithm’s performance is thus significantly faster as compared to existing POMDP solvers which would be exponential in the number of states and thus doubly exponential in the number of components Silver & Veness (2010), Pineau et al.
+(2003).
 If the problem is approached directly as a single POMDP, it will have a prohibitively vast state space of approximately $10^{2000}$ states.
@@
 (2023) demonstrated that standard methods indeed become computationally intractable after a few components due to this combinatorial explosion.
-Figure 6: Log-log plot of computational complexity of the proposed approach for varying numbers of components.
+5.2 Implementation and Evaluation for Financial Loss-Budget Management Scenario
+Our second experimental scenario addresses a portfolio risk management task.
+We use daily price data for the S&P 500 constituent stocks over a two-year window, reserving the final $T=120$ trading days for evaluation and using earlier data for training.
+The core component of the POMDP is an unobserved latent state defined per component as a debit-only loss budget (health), $s_{t}\in[0,100]$.
+Each component receives a small loss budget: on a day with a negative return, $s_{t}$ is debited proportionally and decreases; on a non-negative day, $s_{t}$ is unchanged; the state does not self-recover.
+Health increases only when the agent executes recapitalize.
+All actions draw from one shared, limited budget, and actions are taken when drift relative to the per-component no-loss floor becomes meaningful.
+This design is practice-inspired for two reasons.
+First, because we manage a large number of components, governance and our own policy favor a conservative stance: we avoid repeatedly allocating budget to components with recent serial losses, so the health is debit-only and does not auto-replenish.
+Second, it follows the risk-budgeting workflow described in Benham & Bebee (2024)—set a budget ex ante, allocate and monitor against a benchmark, and treat material drift as a trigger for action.
+To make the benchmark operational, we instantiate a per-component no-loss floor: losses are deviations that consume the per-component budget; gains are consistent with the floor and do not raise limits by themselves; replenishment occurs only through recapitalize.
+For training and evaluation, components are assumed independent.
+Actions and Costs: The agent’s action space $\mathcal{A}=\{\text{defer, inspect, recapitalize}\}$ manages the per-component loss budget (health).
+All actions draw from a shared, limited budget $B$ and follow a strict cost hierarchy $c_{\mathrm{recapitalize}}>c_{\mathrm{inspect}}>c_{\mathrm{defer}}$:
+  • Defer: Continue with the current position. Incurs a low, continuous cost $c_{\mathrm{defer}}$ each step. Health remains subject to depletion by negative returns.
+  • Inspect: Pay $c_{\mathrm{inspect}}$ to obtain a precise observation of the hidden health $s_{t}$ for the selected component.
+  • Recapitalize: Pay the high cost $c_{\mathrm{recapitalize}}$ to rebuild health by resetting $s_{t}$ to $100$. This is the only action that increases health.
+Objective and Failure Condition.
+The agent’s objective is to learn a policy $\pi$ that maximizes its survival time.
+An absorbing failure state is triggered immediately if any component’s health is exhausted, i.e., $s_{t}\leq 0$.
+For each day the agent survives, it receives a reward of +1.
+This setup forces the agent to learn a sophisticated policy that balances the continuous drain from defer costs and market losses against the high, discrete costs of inspection and recapitalization, in order to prolong its survival.
+5.2.1 Analysis of Recapitalization Policy
+We evaluate our approach on a stock-level loss-budget management task constructed from the S&P 500 universe.
+Starting from $500$ constituents, we retain the subset with at least $80\%$ daily-price coverage over the preceding three years, yielding $471$ components.
+As in the infrastructure experiment, we reserve the final $T=120$ trading days for evaluation and use earlier data for model training.
+State, actions, and costs.
+Each component $j$ is modeled as a single-component monotonic POMDP with an unobserved, debit-only loss-budget (health) $s_{t}^{j}\in[0,100]$.
+Negative returns debit $s_{t}^{j}$ proportionally; non-negative returns leave $s_{t}^{j}$ unchanged; the state does not self-recover.
+The action set is $\mathcal{A}=\{\text{defer},\text{inspect},\text{recapitalize}\}$ with a strict cost hierarchy $c_{\mathrm{recapitalize}}>c_{\mathrm{inspect}}>c_{\mathrm{defer}}$.
+A global budget $B_{\text{tot}}$ is shared across all components.
+Table  3 presents the values of the various parameters used for the experiments.
+Parameter Value
+Total budget $B_{\text{tot}}$ 15,00015{,}000
+Recapitalization cost $c_{\textsc{recap}}$ $10.0$
+Inspection cost $c_{\textsc{insp}}$ $0.5$
+Defer cost $c_{\textsc{def}}$ $0.2$
+Number of components $471$
+Evaluation horizon $T$ $120$ days
+Table 3: Cost and budget settings for the stock-level scenario.
+Policies compared.
+We compare four policies:
+  1. 1.
+    Oracle: full observability of the health $s_{t}$; recapitalize whenever $s_{t}<20$ (no inspection cost).
+  2. 2.
+    Oracle-guided meta-PPO: the agent chooses inspect vs.
+defer; upon defer, it executes the oracle’s suggested restorative/default control; upon inspect, it pays $c_{\mathrm{inspect}}$ to reduce belief uncertainty.
+The agent learns when to buy observations and when to accept uncertainty.
+  3. 3.
+    Baseline (Heuristic): fixed inspect every 55 trading days; if the observed $s_{t}<20$, take recapitalize; otherwise defer.
+  4. 4.
+    Vanilla meta-PPO: trained on the same component–budget pairs as the oracle-guided agent but without oracle shaping.
+Budget allocation.
+We allocate $B_{\text{tot}}$ across the $471$ components using the same random-forest surrogate procedure as in the maintenance experiment: for each component $i$ we fit a concave surrogate for the map $B\mapsto\mathbb{E}[T^{i}_{\max}(B)]$ and solve a tractable concave maximization to obtain per-component budgets.
+Training details.
+Both oracle-guided meta-PPO and vanilla meta-PPO are trained for $2\times 10^{6}$ timesteps with Adam step size $10^{-4}$, minibatch size $128$, PPO horizon $T_{\mathrm{PPO}}=2048$, and discount factor $0.95$.
+For each component and policy we run $100$ simulations and report the component-level average $T^{i}_{\max}$; we then average across all $471$ components to obtain $\hat{T}_{\max}$.
+Figure 7: S&P 500 stock-level scenario: average survival time $\hat{T}_{\max}$ under a shared budget $B_{\mathrm{tot}}=15{,}000$ across 471 components.
+Observed ordering: Oracle >> Oracle-guided meta-PPO >> Baseline >> Vanilla meta-PPO.
+Findings.
+Under the shared budget $B_{\text{tot}}=15{,}000$ across $471$ components and a $120$-day evaluation horizon, we observe a consistent ordering (see Figure 7): $\textbf{Oracle}>\textbf{Oracle-guided meta-PPO}>\textbf{Baseline}>\textbf{Vanilla meta-PPO}$.
+The Vanilla meta-PPO tends to conserve budget and rarely recapitalizes, yielding the lowest survival time.
+The Baseline performs periodic inspections (every 55 trading days) and recapitalizes below the threshold but spends budget indiscriminately and misses urgent cases.
+By contrast, the Oracle-guided meta-PPO learns when to inspect versus defer and when to act, allocating budget to higher-value opportunities; it reliably outperforms the Baseline and closes a substantial portion of the gap to the Oracle upper bound.
+5.2.2 Generalizability and Window Robustness of the Oracle-guided Meta-PPO
+Design.
+We vary the number of components $N\in\{5,10,20,100,471\}$.
+For each NN, the policy is trained on rolling $120$-day train windows and evaluated both in-sample (train) and on a held-out test window.
+We report average survival time (days) over r=5r{=}5 seeds; error bars denote $\pm 1$ standard deviation across seeds.
+Figure 8: Oracle-guided meta-PPO: train vs. test across component set size NN. The $y$-axis is average survival time (days); the $x$-axis is the number of components.
+Findings.
+(1) Both train and test curves decrease as NN grows, reflecting budget dilution and increased problem complexity.
+(2) Train performance is consistently above test with a modest generalization gap that tends to widen at larger NN.
+(3) Variability is non-negligible and generally larger at higher NN.
+Takeaway.
+As can be seen from Figure 8, the oracle-guided meta-PPO exhibits window robustness: trends are consistent across train windows, and the train-to-test drop remains moderate.
+At small NN, the effective exploration/interaction budget is limited, which can hinder learning; as NN increases, richer allocation opportunities make better use of the oracle guidance even though absolute survival time declines under a fixed total budget.
 6 Conclusions
-In this paper we introduce a framework to provide computationally efficient solutions, scalable for large scale budget-constrained multi-component monotonic POMDPs.
-Noting that the coupling is expressed solely through the shared budget, our method functions by first decomposing the large $n$-component POMDP into $n$ independent single-component POMDPs using a random forest budget allocation of the shared budget.
-Next, we compute the oracle policy using value iteration for each component POMDP and budget pair.
-Finally, using this oracle policy together with a meta-trained PPO agent, we solve each component POMDP.
-Consequently, we obtain the approximately optimal policy for the multi-component POMDP.
-The experimental evaluations conducted for a real-world scenario of large scale infrastructure maintenance show that the proposed approach outperforms existing baselines and almost matches the performance of the oracle policy, which relies on information not available to the planner in a partially observable setting.
-Furthermore, computational complexity analysis demonstrates the scalability of our framework.
+We proposed a scalable framework for solving budget-constrained multi-component monotonic POMDPs.
+Our chief theoretical contribution is a proof that the single-component value function is concave in budget, which underpins an efficient two-step solution strategy.
+First, a random-forest surrogate exploits that concavity to distribute the shared budget across components, thereby decomposing the large $n$-component POMDP into $n$ independent single-component POMDPs.
+Second, an oracle-guided, meta-trained PPO agent—shaped by value iteration on the fully observable counterpart—learns a near-optimal policy for each component–budget pair.
+Comprehensive experiments on two disparate domains confirm the framework’s generality.
+For a 1000-component building-maintenance task, our method significantly prolongs component survival relative to baseline heuristics and approaches the performance of the oracle policy.
+On an ETF portfolio-rebalancing problem with draw-down–risk budgets, the same algorithm consistently preserves portfolio viability and outperforms vanilla PPO and the equal-weight baseline.
+Across both settings, empirical runtimes grow linearly with the number of components, validating the scalability predicted by our complexity analysis.
 Future work will focus on extending the framework’s capabilities to more dynamic budget allocation schemes and more complicated hierarchical budget constraints.
+Appendix A Function Approximation of $\mathbb{E}[T_{\max}]$
+We model $\mathbb{E}[T^{i}_{\max}]$ as an exponential function of the budget allocated to component ii.
+The choice of an exponential function is motivated by its ability to capture the saturation in $\mathbb{E}[T^{i}_{\max}]$ values at higher budget levels, a result of the finite planning horizon $H$.
+Additionally, the exponential model accounts for non-zero $\mathbb{E}[T^{i}_{\max}]$ even when the budget is zero.
+Figure 9: Exponential $\tilde{T}^{i}_{\max}$ curves obtained using non-linear least-squares regression.
+To validate the accuracy of this exponential model for $\mathbb{E}[T^{i}_{\max}]$, we conducted non-linear least squares regression on 100 infrastructure components.
+Figure 9 illustrates the curves obtained through this regression, where $\mathbb{E}[T^{i}_{\max}]$ is modeled as an exponential function.
+The results indicate that the exponential function provides a strong approximation for $\mathbb{E}[T^{i}_{\max}]$, with an average coefficient of determination $R^{2}_{mean}=0.899$.

### v3

@@
 The primary objective of this paper is to determine a policy $\pi^{*}$ for this multi-component monotonic POMDP over a horizon $H$, that maximizes the sum of expectations of the individual times before reaching the absorbing state for each component, while adhering to the total budget $B$.
-We denote this maximal time $k$ by $T_{max}=\sum_{i=1}^{n}T^{i}_{max}$, where $T^{i}_{max}$ denotes the corresponding maximal time for component ii.
+We denote this maximal time $k$ by $T_{max}=\sum_{i=1}^{n}T^{i}_{max}$, where $T^{i}_{max}$ denotes the corresponding maximal time for component $i$.
 Mathematically, the problem can be formulated as:
@@
 We consider a POMDP defined by the tuple $\langle S,A,T,R,\Omega,O,\gamma,\mathcal{C}\rangle$, where:
-  • $S$ is a finite state space and AA is a finite action space, as defined in Section 3.
+  • $S$ is a finite state space and $A$ is a finite action space, as defined in Section 3.
   • $T(s^{\prime}|s,a)$ is the transition kernel (1); $R(s,a)$ is the reward function.
@@
 Lemma 1 (Concavity under Affine Shift).
-If $f(x)$ is concave on an interval II, then f(xl)f(x-l) is concave on the interval $\{y\mid y=x+l,x\in I\}$ for any constant $l$.
+If $f(x)$ is concave on an interval $I$, then f(xl)f(x-l) is concave on the interval $\{y\mid y=x+l,x\in I\}$ for any constant $l$.
 Proof.
@@
 Proof.
-Under reward scheme (4) each non-absorbing step contributes exactly 11 to the return; steps in state 0 contribute 0. Hence, for any budget-feasible policy $\pi$,
+Under reward scheme (4) each non-absorbing step contributes exactly $1$ to the return; steps in state 0 contribute 0. Hence, for any budget-feasible policy $\pi$,
 $\text{Total reward}=\mathbb{E}\Bigl{[}\sum_{t=0}^{H}\mathbf{1}\{s_{t}\neq 0\}\Bigr{]}=\mathbb{E}[T_{\max}(\pi)].$
@@
 We now integrate the approaches described in Section 4.2 and Section 4.3 to compute the optimal policy for an $n$-component POMDP, where $n$ is substantially large.
-Utilizing the random forest regressor, we efficiently approximate $\mathbb{E}[T_{\max}]$ for each component ii.
+Utilizing the random forest regressor, we efficiently approximate $\mathbb{E}[T_{\max}]$ for each component $i$.
 Additionally, we meta-train the oracle-guided PPO agent by continuously updating the policy network’s parameters over a randomly selected subset of components and budget values.
@@
 These pairs correspond to the run-averaged $T^{i}_{\max}$ values and the respective budget values $b^{i}$ from Section 5.1.1.
-The input to the random forest model is a vector consisting of the shape and scale factors of the Weibull distribution, which represent $\mathbb{E}[T]$ and $\sigma^{2}_{\mathbb{E}[T]}$, along with the replacement and inspection costs for a given component ii.
+The input to the random forest model is a vector consisting of the shape and scale factors of the Weibull distribution, which represent $\mathbb{E}[T]$ and $\sigma^{2}_{\mathbb{E}[T]}$, along with the replacement and inspection costs for a given component $i$.
 If a different distribution was used to model the transition probability, we would similarly extract the parameters, $\mathbb{E}[T]$ and $\sigma^{2}_{\mathbb{E}[T]}$, for inclusion in the input vector.
@@
 5.1.3 Analysis of Time Complexity
-Finally, we analyze the time complexity of our proposed approach for varying number of components NN. As mentioned earlier, our method comprises of four major steps:
+Finally, we analyze the time complexity of our proposed approach for varying number of components $N$. As mentioned earlier, our method comprises of four major steps:
   1. 1.
-    Random Forest regression for estimating $\tilde{T}^{i}_{max}$ for each component ii.
+    Random Forest regression for estimating $\tilde{T}^{i}_{max}$ for each component $i$.
   2. 2.
@@
   3. 3.
-    Baseline (Heuristic): fixed inspect every 55 trading days; if the observed $s_{t}<20$, take recapitalize; otherwise defer.
+    Baseline (Heuristic): fixed inspect every $5$ trading days; if the observed $s_{t}<20$, take recapitalize; otherwise defer.
   4. 4.
@@
 The Vanilla meta-PPO tends to conserve budget and rarely recapitalizes, yielding the lowest survival time.
-The Baseline performs periodic inspections (every 55 trading days) and recapitalizes below the threshold but spends budget indiscriminately and misses urgent cases.
+The Baseline performs periodic inspections (every $5$ trading days) and recapitalizes below the threshold but spends budget indiscriminately and misses urgent cases.
 By contrast, the Oracle-guided meta-PPO learns when to inspect versus defer and when to act, allocating budget to higher-value opportunities; it reliably outperforms the Baseline and closes a substantial portion of the gap to the Oracle upper bound.
@@
 We vary the number of components $N\in\{5,10,20,100,471\}$.
-For each NN, the policy is trained on rolling $120$-day train windows and evaluated both in-sample (train) and on a held-out test window.
+For each $N$, the policy is trained on rolling $120$-day train windows and evaluated both in-sample (train) and on a held-out test window.
 We report average survival time (days) over r=5r{=}5 seeds; error bars denote $\pm 1$ standard deviation across seeds.
@@
 Findings.
-(1) Both train and test curves decrease as NN grows, reflecting budget dilution and increased problem complexity.
-(2) Train performance is consistently above test with a modest generalization gap that tends to widen at larger NN.
-(3) Variability is non-negligible and generally larger at higher NN.
+(1) Both train and test curves decrease as $N$ grows, reflecting budget dilution and increased problem complexity.
+(2) Train performance is consistently above test with a modest generalization gap that tends to widen at larger $N$.
+(3) Variability is non-negligible and generally larger at higher $N$.
 Takeaway.
 As can be seen from Figure 8, the oracle-guided meta-PPO exhibits window robustness: trends are consistent across train windows, and the train-to-test drop remains moderate.
-At small NN, the effective exploration/interaction budget is limited, which can hinder learning; as NN increases, richer allocation opportunities make better use of the oracle guidance even though absolute survival time declines under a fixed total budget.
+At small $N$, the effective exploration/interaction budget is limited, which can hinder learning; as $N$ increases, richer allocation opportunities make better use of the oracle guidance even though absolute survival time declines under a fixed total budget.
 6 Conclusions
@@
 Appendix A Function Approximation of $\mathbb{E}[T_{\max}]$
-We model $\mathbb{E}[T^{i}_{\max}]$ as an exponential function of the budget allocated to component ii.
+We model $\mathbb{E}[T^{i}_{\max}]$ as an exponential function of the budget allocated to component $i$.
 The choice of an exponential function is motivated by its ability to capture the saturation in $\mathbb{E}[T^{i}_{\max}]$ values at higher budget levels, a result of the finite planning horizon $H$.